# Guided Flow Policy: Learning from High-Value Actions in Offline Reinforcement Learning

### v1

@@ Appendix A Implementation details
Doubled Q-learning. Following standard practice, two separate critic networks are trained and then aggregated to compute action values, either by taking the mean or minimum fujimoto2018addressing. As detailed in Sec. 4.3, we find that the aggregation function has a significant impact on performance for specific tasks. Specifically, by reevaluating ReBRAC on OGBench using the same aggregation function as GFP and FQL, we achieved substantial performance improvements.
Minibatch size. We use a minibatch size of $B=256$ across most experiments, except on the most challenging tasks, where we evaluate each method with both $B=256$ and $B=1024$. The humanoidmaze-large-navigate environment is the only task where methods benefit from different batch sizes: GFP performs best with $B=1024$, while other methods work better with $B=256$. Note that on this task, using $\gamma=0.999$ substantially improves the performance of ReBRAC and FQL compared to previously reported results. For Minari Gym-Mujoco, we use $B=1024$ following the recommendation in tarasov2023revisiting for D4RL Gym-Mujoco.
Training and evaluation. To ensure a fair comparison with FQL, we use identical training durations: 1 M gradient steps on OGBench state-based environments, and 500 K steps on D4RL and OGBench pixel-based tasks. For Minari, we adopt 1 M steps following standard practices. Evaluation differs according to the benchmark: D4RL and Minari scores are computed at the end of training, while OGBench scores are averaged over the final three checkpoints (800K, 900K, and 1M steps for state-based, 300K, 400K and 500K steps for pixel-based) following their official evaluation protocol. All results are reported across 8 random seeds that were not used during the hyperparameter tuning process, and each evaluation is done over 100 episodes. Except for OGBench pixel-based tasks, where scores are computed using 50 episodes and 4 seeds, due to high computational cost, following park2025flow.
Minari MuJoCo reference scores. Since reference scores for the MuJoCo locomotion tasks are not yet available in Minari minari, we use the reference scores reported in its predecessor D4RL fu2020d4rl.
@@ B.2 Advantage weight similar to AWR
As explained at the end of Sec. 3, our $g_{\eta}$ is a soft-max weight comparing $a$ to $a^{\pi_{\theta}}$, an action sampled from the actor, but one could alternatively use an advantage weighted term similar to AWR peng2019advantage, but using the actor to compute a baseline:
@@ B.5 Sensitivity analysis to $\eta$ and $\alpha$
To study how sensitive GFP is to the temperature $\eta$ and the BC coefficient $\alpha$, we evaluated 8 variations around task-specific hyperparameters $(\alpha^{*},\eta^{*})$ on 12 tasks, with comparison to FQL sensitivity to $\alpha$. These experiments reveal that $\alpha$ is the most important hyperparameter of GFP, as it is for most offline RL methods tarasov2023revisiting; park2024value, and that GFP is less sensitive to precise $\eta$ tuning. It justifies our methodology for hyperparameter search, by first fixing $\eta=10^{-3}$, and then sweeping over $\eta$ only once $\alpha$ is chosen. From the results averaged on 12 tasks, it can be observed that $\alpha$ and $\eta$ are partially correlated: when increasing $\alpha$ (i.e. stronger regularization), it is better to decrease $\eta$ (stronger filtering).
@@
Figure 7: Sensitivity analysis to the BC coefficient $\alpha$, and to the temperature $\eta$, in log-scale. Evaluates the sensitivity by testing variations around task-specific hyperparameters $(\alpha^{*},\eta^{*})$ reported Tab. 13, using 8 seeds. Plots (a), (c) and (e), report GFP’s sensitivity to $\alpha^{*}$ and $\eta^{*}$, on antmaze-large-navigate-task1, cube-double-play-task2, and scene-play-task2, respectively. These three tasks were chosen following FQL park2025flow ablation study on $\alpha$ (FQL page 15). Plots (b), (d) and (f), report the sensitivity of FQL with respect to $\alpha$ (reevaluated as FQL’s authors did not share exact numbers). Plots (g) and (h) aggregate the analysis over 12 tasks (cube single and double, play and noisy; scene play and noisy; antmaze navigate large and giant; humanoidmaze medium navigate and stitch, where for each environment we use the default task). These experiments reveal that $\alpha$ is the most important hyperparameter of GFP, as it is for most offline RL methods tarasov2023revisiting; park2024value, and that GFP is less sensitive to precise $\eta$ tuning.
@@ Appendix C Complete results over 144 tasks
Table 8: Offline RL full results on OGBench, part 1/2. Models were trained with 8 random seeds and evaluated over 100 episodes, following the setup of prior work park2024ogbench; park2025flow. Scores are averaged across seeds; values within 95% of the best performance are shown in bold, while italics indicate scores reported from prior work park2025flow. GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training. The $\pm$ symbol denotes the standard deviation over seeds. See Tab 9 for the other tasks.
@@
Table 9: Offline RL full results on OGBench, part 2/2. Models were trained with 8 random seeds and evaluated over 100 episodes, following the setup of prior work park2024ogbench; park2025flow. Scores are averaged across seeds; values within 95% of the best performance are shown in bold, while italics indicate scores reported from prior work park2025flow. GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training. The $\pm$ symbol denotes the standard deviation over seeds. See Tab 8 for the other tasks.
@@
1 Following park2025flow to reduce the computational cost of pixel-based tasks, for these tasks we use 4 seeds and 50 episodes per seed; moreover, we chose not to evaluate the VaBC policy, as it is only a byproduct.
Table 10: Offline RL Results on OGBench Tasks (only task evaluated in FQL): Performance comparison across different offline RL algorithms. All prior work results are taken from park2025flow to compare a wide range of previous methods against GFP.
@@
Table 11: Offline RL full results on D4RL. For each task, models were trained with 8 random seeds and evaluated at the end of training. Reported values are the average normalized scores over the final 100 evaluation episodes, with $\pm$ denoting the standard deviation across seeds. Italics indicate scores from prior work fu2020d4rl; fujimoto2021minimalist; tarasov2023revisiting; park2025flow, and bold denotes values within 95% of the best performance. GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training.

### v2

@@ Appendix A Implementation details
Doubled Q-learning. Following standard practice, two separate critic networks are trained and then aggregated to compute action values, either by taking the mean or minimum (Fujimoto et al., 2018). As detailed in Sec. 4.3, we find that the aggregation function has a significant impact on performance for specific tasks. Specifically, by reevaluating ReBRAC on OGBench using the same aggregation function as GFP and FQL, we achieved substantial performance improvements.
Minibatch size. We use a minibatch size of $B=256$ across most experiments, except on the most challenging tasks, where we evaluate each method with both $B=256$ and $B=1024$. The humanoidmaze-large-navigate environment is the only task where methods benefit from different batch sizes: GFP performs best with $B=1024$, while other methods work better with $B=256$. Note that on this task, using $\gamma=0.999$ substantially improves the performance of ReBRAC and FQL compared to previously reported results. For Minari Gym-Mujoco, we use $B=1024$ following the recommendation in Tarasov et al. (2023) for D4RL Gym-Mujoco.
Training and evaluation. To ensure a fair comparison with FQL, we use identical training durations: 1 M gradient steps on OGBench state-based environments, and 500 K steps on D4RL and OGBench pixel-based tasks. For Minari, we adopt 1 M steps following standard practices. Evaluation differs according to the benchmark: D4RL and Minari scores are computed at the end of training, while OGBench scores are averaged over the final three checkpoints (800K, 900K, and 1M steps for state-based, 300K, 400K and 500K steps for pixel-based) following their official evaluation protocol. All results are reported across 8 random seeds that were not used during the hyperparameter tuning process, and each evaluation is done over 100 episodes. Except for OGBench pixel-based tasks, where scores are computed using 50 episodes and 4 seeds, due to high computational cost, following Park et al. (2025).
Minari MuJoCo reference scores. Since reference scores for the MuJoCo locomotion tasks are not yet available in Minari (Younis et al., 2024), we use the reference scores reported in its predecessor D4RL (Fu et al., 2020).
@@ B.1 Modified Bellman target
Implementation Note on Eq. 7: After finishing the experiments, we realized that in our implementation of the modified $y^{\text{VaBC}}$ Bellman target, we were computing $Q_{\bar{\phi}}(s^{\prime},\mu_{\omega}(s,z))$ instead of $Q_{\bar{\phi}}(s^{\prime},\mu_{\omega}(s^{\prime},z))$. In our public release of the code, we kept this behavior by default to match the code used in our experiments, but we added an option to use $Q_{\bar{\phi}}(s^{\prime},\mu_{\omega}(s^{\prime},z))$, for future experiments.
@@ B.2 Advantage weight similar to AWR
As explained at the end of Sec. 3, our $g_{\eta}$ is a soft-max weight comparing $a$ to $a^{\pi_{\theta}}$, an action sampled from the actor, but one could alternatively use an advantage weighted term similar to AWR (Peng et al., 2019), but using the actor to compute a baseline:
@@ B.5 Sensitivity analysis to $\eta$ and $\alpha$
To study how sensitive GFP is to the temperature $\eta$ and the BC coefficient $\alpha$, we evaluated 8 variations around task-specific hyperparameters $(\alpha^{*},\eta^{*})$ on 12 tasks, with comparison to FQL sensitivity to $\alpha$. These experiments reveal that $\alpha$ is the most important hyperparameter of GFP, as it is for most offline RL methods (Tarasov et al., 2023; Park et al., 2024b), and that GFP is less sensitive to precise $\eta$ tuning. It justifies our methodology for hyperparameter search, by first fixing $\eta=10^{-3}$, and then sweeping over $\eta$ only once $\alpha$ is chosen. From the results averaged on 12 tasks, it can be observed that $\alpha$ and $\eta$ are partially correlated: when increasing $\alpha$ (i.e. stronger regularization), it is better to decrease $\eta$ (stronger filtering).
@@
Figure 7: Sensitivity analysis to the BC coefficient $\alpha$, and to the temperature $\eta$, in log-scale. Evaluates the sensitivity by testing variations around task-specific hyperparameters $(\alpha^{*},\eta^{*})$ reported Tab. 13, using 8 seeds. Plots (a), (c) and (e), report GFP’s sensitivity to $\alpha^{*}$ and $\eta^{*}$, on antmaze-large-navigate-task1, cube-double-play-task2, and scene-play-task2, respectively. These three tasks were chosen following FQL (Park et al., 2025) ablation study on $\alpha$ (FQL page 15). Plots (b), (d) and (f), report the sensitivity of FQL with respect to $\alpha$ (reevaluated as FQL’s authors did not share exact numbers). Plots (g) and (h) aggregate the analysis over 12 tasks (cube single and double, play and noisy; scene play and noisy; antmaze navigate large and giant; humanoidmaze medium navigate and stitch, where for each environment we use the default task). These experiments reveal that $\alpha$ is the most important hyperparameter of GFP, as it is for most offline RL methods (Tarasov et al., 2023; Park et al., 2024b), and that GFP is less sensitive to precise $\eta$ tuning.
@@ Appendix C Complete results over 144 tasks
Table 8: Offline RL full results on OGBench, part 1/2. Models were trained with 8 random seeds and evaluated over 100 episodes, following the setup of prior work (Park et al., 2024a; 2025). Scores are averaged across seeds; values within 95% of the best performance are shown in bold, while italics indicate scores reported from prior work (Park et al., 2025). GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training. The $\pm$ symbol denotes the standard deviation over seeds. See Tab 9 for the other tasks.
@@
Table 9: Offline RL full results on OGBench, part 2/2. Models were trained with 8 random seeds and evaluated over 100 episodes, following the setup of prior work (Park et al., 2024a; 2025). Scores are averaged across seeds; values within 95% of the best performance are shown in bold, while italics indicate scores reported from prior work (Park et al., 2025). GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training. The $\pm$ symbol denotes the standard deviation over seeds. See Tab 8 for the other tasks.
@@
1 Following Park et al. (2025) to reduce the computational cost of pixel-based tasks, for these tasks we use 4 seeds and 50 episodes per seed; moreover, we chose not to evaluate the VaBC policy, as it is only a byproduct.
Table 10: Offline RL Results on OGBench Tasks (only task evaluated in FQL): Performance comparison across different offline RL algorithms. All prior work results are taken from Park et al. (2025) to compare a wide range of previous methods against GFP.
@@
Table 11: Offline RL full results on D4RL. For each task, models were trained with 8 random seeds and evaluated at the end of training. Reported values are the average normalized scores over the final 100 evaluation episodes, with $\pm$ denoting the standard deviation across seeds. Italics indicate scores from prior work (Fu et al., 2020; Fujimoto and Gu, 2021; Tarasov et al., 2023; Park et al., 2025), and bold denotes values within 95% of the best performance. GFP actor $\pi_{\theta}$ is our primary policy, while GFP VaBC $\pi_{\omega}$ is reported as a byproduct of training.